# Algebraic non-Hermitian skin effect and unified non-Bloch band theory in arbitrary dimensions

### v1

@@ Appendix B The GBZ formula in three-dimensional systems
We set the Hamiltonian parameters to $t_{x}=t_{y}=1/2$, $t_{z}=1$, $t_{xyz}=i/2$, and take the system size as $L_{x}=L_{y}=L_{z}=L=12$. The corresponding OBC eigenvalues are plotted by the red dots in Fig. 5(a). We pick up a generic OBC eigenvalue $E_{0}=1.55391-0.22258i$ (the blue dot) and calculate its GFS. According to our formula discussed in Sec. V, the bulk allowable $\rho_{z}$ values can be solved as the eigenvalues of the transfer matrix $\mathbb{T}(E_{0})$. These $2L^{2}$ values of $\rho_{z}$ are plotted by the blue dots in Fig. 5(b), which spans a finite area in the complex $\rho_{z}$ plane. Hence, we have $\dim{\rho_{z}}=2$. For each fixed $\rho_{z}$, the 3D Hamiltonian reduces to a 2D subsystem, which can be solved using the layer transfer matrix approach once again. Here, we show the results that for a fixed $\rho_{z_{0}}=2.15655-1.03812i$, the sub GFS for the 2D subsystem is solved and its projections onto $\rho_{y}$ and $\beta_{x}$ planes are illustrated in Figs. 5(c) and (d), respectively. Here, for the 2D subsystem, layer-$y$ basis has been selected. For each fixed $\rho_{z}$, the GFS for the corresponding 2D subsystem forms 1D curves in ($\beta_{x},\rho_{y}$) space. Meanwhile, $\rho_{z}$ spans a finite region. Consequently, GFS corresponding to a specific energy $E_{0}$ manifests as a 3D manifold. As $E_{0}$ sweeps through the entire OBC continuum spectrum, the collection of GFSs constitutes the GBZ, which is a 5D manifold. Therefore, through Figs. 5(a)-(d), we numerically verified the dimensions for GFS and GBZ using the 3D tight-binding model in Eq. (B).
Now we examine the OBC wavefunctions using the GFS/GBZ basis. According to Eq. (13), the number of standing wave conditions is $12d-\dim{\text{GBZ}}=$ 1. It means that one can enforce the standing wave condition along either one direction. Based on the basis we choose above, the standing waves are formed in the $x$ direction. An illustration of a standing wave component of an OBC eigen wavefunction is shown in Figs. 5(e). The standing wave is first formed along the $x$ direction, and then transfer along the $y$ and $z$ directions. Therefore, the OBC wavefunction can be constructed as follows
@@
where $\rho_{z}$, $\rho_{y}$, and $\beta_{x}$ are taken on the 3D GFS of energy $E_{0}$. The superposition coefficients $A_{ij}$ can be determined by applying the open boundary conditions (or termed zero boundary conditions). We obtain the coefficients $A_{ij}$ by solving the boundary matrix, and plot the constructed wavefunction of Eq. (56) in Figs. 5(f), which fully coincides with the numerical wavefunction by diagonalizing the Hamiltonian matrix. This agreement in OBC wavefunction validates the 3D GBZ formula presented in Sec. V.

### v2

@@ Appendix G The GFS formula in three-dimensional systems
We set the Hamiltonian parameters to $t_{x}=t_{y}=1/2$, $t_{z}=1$, $t_{xyz}=i/2$, and take the system size as $L_{x}=L_{y}=L_{z}=L=12$. The corresponding OBC eigenvalues are plotted by the red dots in Fig. 13(a). We pick up a generic OBC eigenvalue $E_{0}=1.55391-0.22258i$ (the blue dot) and calculate its GFS. According to our formula discussed in Sec. V, the bulk allowable $\rho_{z}$ values can be solved as the eigenvalues of the transfer matrix $\mathbb{T}(E_{0})$. These $2L^{2}$ values of $\rho_{z}$ are plotted by the blue dots in Fig. 13(b), which spans a finite area in the complex $\rho_{z}$ plane. Hence, we have $\dim{\rho_{z}}=2$. For each fixed $\rho_{z}$, the 3D Hamiltonian reduces to a 2D subsystem, which can be solved using the layer transfer matrix approach once again. Here, we show the results that for a fixed $\rho_{z_{0}}=2.15655-1.03812i$, the sub GFS for the 2D subsystem is solved and its projections onto $\rho_{y}$ and $\beta_{x}$ planes are illustrated in Figs. 13(c) and (d), respectively. Here, for the 2D subsystem, layer-$y$ basis has been selected. For each fixed $\rho_{z}$, the GFS for the corresponding 2D subsystem forms 1D curves in ($\beta_{x},\rho_{y}$) space. Meanwhile, $\rho_{z}$ spans a finite region. Consequently, GFS corresponding to a specific energy $E_{0}$ manifests as a 3D manifold. As $E_{0}$ sweeps through the entire OBC continuum spectrum, the collection of GFSs constitutes the GBZ, which is a 5D manifold. Therefore, through Figs. 13(a)-(d), we numerically verified the dimensions for GFS and GBZ using the 3D tight-binding model in Eq. (G).
Now we examine the OBC wavefunctions using the GFS basis. According to Eq. (13), the number of standing wave conditions is $2d-\dim{\text{GBZ}}=1$. It means that one can enforce the standing wave condition along either one direction. Based on the basis we choose above, the standing waves are formed in the $x$ direction. An illustration of a standing wave component of an OBC eigen wavefunction is shown in Figs. 13(e). The standing wave is first formed along the $x$ direction, and then transfer along the $y$ and $z$ directions. Therefore, the OBC wavefunction can be constructed as follows
@@
where $\rho_{z}$, $\rho_{y}$, and $\beta_{x}$ are taken on the 3D GFS of energy $E_{0}$. The superposition coefficients $A_{ij}$ can be determined by applying the open boundary conditions (or termed zero boundary conditions). We obtain the coefficients $A_{ij}$ by solving the boundary matrix, and plot the constructed wavefunction of Eq. (87) in Figs. 13(f), which fully coincides with the numerical wavefunction by diagonalizing the Hamiltonian matrix. This agreement in OBC wavefunction validates the 3D GFS formula presented in Sec. V.